# Evaluating Sparse Autoencoders for Monosemantic Representation

### v1

@@ Introduction
Moreover, Range Masking, Adaptive Dampening, and Full Masking each rely on a saliency threshold, which is a hyperparameter that selects the top-$p\%$ of neurons deemed most relevant to the target concept. We set $p=0.1$ for IMDB, $p=0.3$ for AG News, and $p=0.2$ for the remaining datasets. In contrast, AURA (Suau et al. 2024) and our method, APP, do not require this hyperparameter. Both apply interventions across all neurons in the selected layer, avoiding the need to tune or justify a saliency cutoff. However, because SAE neurons typically activate on only a small subset of inputs for any given concept, we introduce an activation-frequency threshold $\tau$ to ensure reliability. Specifically, for each SAE neuron $h^{l}_{j}$, and each concept $c_{i}$, let $\mathcal{X}_{c_{i}}$ be the set of corresponding input samples. We define the firing frequency of $h^{l}_{j}$ on $c_{i}$ as
@@
We apply all interventions at a consistent computation point, immediately before the language-model head, ensuring comparability across models and methods. Specifically, we intervene after the MLP and residual-stream addition in the final transformer block of the model. For SAEs, we apply interventions just after the SAE’s activation nonlinearity.
To isolate the causal impact on prediction, interventions are restricted to the token corresponding to the model’s output label. This allows us to precisely assess the influence of each concept-erasure method on the final decision, without confounding effects from earlier tokens.
Results and Analysis
Comparison of Intervention Effectiveness: SAEs vs. Base Model. As it can be seen in Table 2, across nearly all settings, we find that partial intervention methods (particularly APP and Aura) consistently achieve higher $\Delta_{\mathrm{Acc}}$ and $\Delta_{\mathrm{Conf}}$ when applied to SAE representations compared to the base model. Specifically, for APP, SAE-based interventions outperformed the base model in 36 out of 40 comparisons. For Aura, SAE-based interventions were more effective in all 40 cases. In contrast, full masking shows little benefit from SAE representations: only 11 out of 40 interventions resulted in better outcomes than when applied to the base model. This discrepancy suggests that coarse suppression methods fail to capitalize on the increased concept separability offered by SAEs. These findings reinforce that SAE representations are more disentangled and that fine-grained, distribution-aware methods are better equipped to exploit this structure for effective concept removal. Furthermore, focusing specifically on the APP intervention (which is a partial intervention), we observe that within the SAE family, increasing capacity consistently enhances intervention quality. In particular, $\Delta_{\mathrm{Conf}}$ consistently increases as we scale from 16 k to 65 k latent dimensions, reflecting improved confidence suppression for the target concept. $\Delta_{\mathrm{Acc}}$ also improves across most datasets, with only minor exceptions, further underscoring the role of latent dimensionality in enabling more precise and effective concept removal.
Partial Interventions Vs Full Interventions. Full masking ranks as the worst‐performing method in 36 out of 40 $\Delta$-metrics on the SAEs and in 1 out of 10 $\Delta$-metrics on the base model. and it also produces the largest perplexity drop in all DPPL evaluations across SAEs and the base model. This underscores that distribution-aware partial methods (e.g., APP, AURA), which leverage activation distributions, are far more effective for targeted concept removal than the coarse, distribution-agnostic full-masking approach.
[…]
Comparative Analysis of APP and AURA. The superior effectiveness of AURA and APP on both $\Delta_{\mathrm{Acc}}$ and $\Delta_{\mathrm{Conf}}$ stems from the fact that they explicitly model not only the target‐concept distribution but also the distributions of all auxiliary concepts. By calibrating their interventions to maximize disruption of $c$ while minimizing collateral effects on $c^{\prime}\neq c$, both methods achieve higher $\Delta$ values than approaches that consider only the target distribution. Between these two, APP pulls ahead of AURA because it leverages fine‐grained, activation‐specific damping rather than a single, per‐neuron factor. AURA mutes an “expert” neuron uniformly, regardless of whether a particular activation is highly characteristic of the target concept, whereas APP computes $\pi_{j,i}(x)$ for each activation and suppresses only the portions of the distribution uniquely associated with $c_{i}$. This activation‐aware attenuation not only sharpens concept removal (higher $\Delta_{\mathrm{Acc}}$ and $\Delta_{\mathrm{Conf}}$) but also preserves the model’s overall fluency more effectively (smaller $\Delta\mathrm{PPL}$).
Related Works
Sparse Autoencoders for Feature Discovery
Sparse Autoencoders (SAEs) have emerged as a powerful method for learning interpretable, monosemantic features from neural network activations (Huben et al. 2024). Recent advances have focused on improving reconstruction quality and scaling through architectural and training strategy innovations such as JumpReLU activations, BatchTopK sparsity, gated, and end-to-end training frameworks (Rajamanoharan et al. 2024a, b; Gao et al. 2025; Bussmann, Leask, and Nanda 2024; Braun et al. 2024). Empirical analyses have validated SAEs’ ability to discover meaningful structures across different domains, from vision-language models to algorithmic patterns like temporal difference learning in LLMs (Sun et al. 2025b; Pach et al. 2025; Demircan et al. 2025). Moreover, evaluation studies have highlighted both their utility for interpretability tasks and remaining challenges with polysemantic representations (Kantamneni et al. 2025; Minegishi et al. 2025). However, none of these empirical analyses evaluated the separability of activation distributions in SAEs as a measure of polysemanticity.
SAE-Based Model Control
The interpretable features learned by SAEs enable precise control over language model behavior. Several works have demonstrated effective steering by carefully selecting and manipulating SAE features, with approaches ranging from supervised methods for identifying relevant dimensions to frameworks using hypernetworks (Arad, Mueller, and Belinkov 2025; He et al. 2025a, b; Bayat et al. 2025; Sun et al. 2025a). However, prior SAE-based control methods did not utilize posterior probabilities, limiting their precision.
Base Model Control and Causal Analysis
Complementing SAE-based approaches, researchers have developed techniques for direct activation control and causal analysis in the base language models. General frameworks for transporting activations facilitate intervention across model architectures (Rodriguez et al. 2025), while causal tracing methods enable precise localization and editing of specific knowledge or biases (Vig et al. 2020; Meng et al. 2022, 2023). These approaches offer foundational tools for probing and manipulating model behavior at the activation level. However, again, they do not leverage posterior probabilities for better intervention and have been applied exclusively to base models, not to representations learned by SAEs.
Conclusion
This work presents the first quantitative analysis of monosemanticity in SAEs compared to their dense base models. We show that while SAEs reduce polysemanticity and improve concept separability, increased sparsity does not always enhance interpretability and often comes with trade-offs in downstream performance. To better characterize monosemanticity, we introduce an activation distribution-aware concept separability score based on the Jensen–Shannon distance, which captures fine-grained distinctions in neuron activations across concepts. We also demonstrate that SAEs support more precise concept-level interventions than base models, particularly when using partial suppression. Building on this, we propose a new method, Attenuation via Posterior Probabilities, which achieves more targeted and effective concept removal with minimal side effects.
Appendix A AG News Neurons Activation
As shown in Figure 4, certain neurons in both the base model and SAEs (e.g., leftmost plots) exhibit considerable overlap in their activation distributions across the four classes, indicating limited class discrimination. In contrast, neurons shown in the middle and right columns reveal more separable activation patterns.
Appendix B The Histogram implementation of KDE
Directly computing the probability densities $f_{h^{l}_{j}\mid c_{i}}(x)$ at inference time is computationally expensive, as it requires comparing each query activation $x$ against all $N$ training activations for every neuron–concept pair. To mitigate this overhead, we adopt a histogram-based kernel density estimation (KDE) approach.
During training, we discretize each activation distribution into $B$ uniform-width histogram bins and store the corresponding bin centers, counts, bandwidths, and normalization constants. At inference time, we compute densities by evaluating each query activation only against these $B$ stored bin centers, avoiding costly comparisons with the full training set. This enables efficient computation using precomputed log-counts and normalization factors, greatly accelerating the density estimation process.
In all experiments, we fix $B=2048$. Let $F$ denote the number of neurons (i.e., distributions), $N$ the number of training samples, and $Q$ the number of query points. Then the computational and memory complexity of our method compared to naïve KDE is:
  • Inference time:
    $\text{Histogram-KDE: }\mathcal{O}(F\,B\,Q),\quad\text{Naïve KDE: }\mathcal{O}(F\,N\,Q)$
  • Memory footprint:
    $\text{Histogram-KDE: }\mathcal{O}(F\,B),\quad\text{Naïve KDE: }\mathcal{O}(F\,N)$
Since $B\ll N$, this substitution yields an approximate $\tfrac{N}{B}$-fold improvement in both inference speed and memory usage. While the approximation may introduce a negligible loss in density accuracy, it provides substantial practical gains in scalability.
Appendix C Details of Experiments
The detailed metrics ($D_{\mathrm{Acc}}$, $D^{\prime}_{\mathrm{Acc}}$, $D_{\mathrm{Conf}}$, and $D^{\prime}_{\mathrm{Conf}}$) before subtraction are reported in Table 3; the corresponding values after subtraction are presented in the main paper (Table 2).
Appendix D Separability Score Trend in Layer 24
As shown in Figure 5, there is a general trend that increasing sparsity tends to improve separability. However, this is not always the case. For example, in subplots (a), (b), and (e), increased sparsity actually leads to reduced separability. Similarly, while performance typically decreases with higher sparsity, there are notable exceptions. In subplots (a) and (e) for instance, performance improves despite the activations becoming more sparse.
@@ Appendix E All SAE Neurons Analysis
While our salient neuron analysis provides valuable insight into the most strongly responding neurons, it considers only a narrow slice of the activation space (specifically, the top 50 neurons per concept). This limited scope may miss neurons that, although not highly ranked by mean activation, are still consistently active across multiple concepts and contribute to polysemanticity. To address this limitation, we broaden our analysis to include all neurons that exhibit non-zero activation for any concept, offering a more comprehensive view of concept overlap beyond the most salient neurons. This extended analysis is conducted only for SAEs, since in the dense base model, all neurons are active across all inputs, rendering such overlap statistics uninformative. The results are shown in Figure 6, which reports the intersection-over-union of active neurons across concepts. By capturing both highly active and more subtly engaged neurons, this analysis reveals a complete picture of polysemantic behavior. Consistent with our earlier findings shown in Figure 1, we observe that higher-capacity SAEs (e.g., 65k dimensions) exhibit lower neuron overlap than their lower-capacity counterparts across all datasets except SST2. Together, Figures 1 and 6 demonstrate that while SAEs significantly reduce polysemanticity, they do not eliminate it entirely. Polysemantic neurons remain present, albeit to a lesser extent.

### v2

@@ Appendix D Implementation Details
Hyperparameters. Range Masking, Adaptive Dampening, and Full Masking each rely on a saliency threshold hyperparameter that determines the top-$p\%$ of neurons considered most relevant to a given concept. For Gemma, we set $p=0.3$ for POS, AG News and NER, $p=0.2$ for Emotions and DBpedia. For DeepSeek, we use $p=0.3$, $0.4$, $0.1$, $0.2$, and $0.4$ for POS, AG News, Emotions, DBpedia, and NER, respectively. In contrast, AURA Suau et al. (2024) and our method, APP, do not require this hyperparameter. Both apply interventions across all neurons in the selected layer, avoiding the need to tune or justify a saliency cutoff. However, because SAE neurons typically activate on only a small subset of inputs for any given concept, we introduce an activation-frequency threshold $\tau$ to ensure reliability. Specifically, for each SAE neuron $h^{l}_{j}$, and each concept $c_{i}$, let $\mathcal{X}_{c_{i}}$ be the set of corresponding input samples. We define the firing frequency of $h^{l}_{j}$ on $c_{i}$ as
@@
Histogram-based KDE. To estimate the densities $f_{h^{l}_{j}\mid c_{i}}(x)$ required by APP, we use a kernel density estimation (KDE) implemented via a histogram-based approximation that preserves accuracy while greatly improving efficiency. During training, each neuron–concept activation distribution is discretized into $B$ uniform-width bins (we use $B=2048$) and we store the bin centers, counts, bandwidths, and normalization constants. At inference, a query activation $x$ is evaluated only against these $B$ centers rather than all $N$ training activations. Let $F$ be the number of neurons and $Q$ the number of query points; this reduces complexity from $\mathcal{O}(FNQ)$ time and $\mathcal{O}(FN)$ memory for naïve KDE to $\mathcal{O}(FBQ)$ and $\mathcal{O}(FB)$, respectively, yielding an approximate $N/B$-fold improvement in both inference speed and memory usage ($B\ll N$).
Intervention Location and Scope. All interventions are applied at a consistent computation point (immediately before the language-model head) to ensure comparability across models and methods. Specifically, interventions are introduced after the MLP and residual addition in the final transformer block. For SAEs, they are applied directly after the SAE activation nonlinearity. To isolate causal effects on prediction, interventions are restricted to the token corresponding to the model’s output label. This design allows precise measurement of each concept-erasure method’s influence on the final decision while avoiding confounding effects from earlier tokens.
Dataset Statistics. We report the dataset sizes and splits used in all experiments to ensure transparency and reproducibility. The AG News dataset contains 120,000 training and 7,600 test examples. The DBpedia dataset includes 100,000 training and 25,000 test examples, both randomly sampled using a fixed seed to ensure reproducibility. For Emotions, we use 16,000 training and 2,000 test examples. The POS Tagging and NER datasets each consist of 100,000 training and 10,000 test examples, also sampled with a fixed random seed.
Model Specifications. As described in the main text, we use the Gemma-2 model with 2 billion parameters (Hugging Face name: google/gemma-2-2b). For DeepSeek, we employ the DeepSeek-R1 model, specifically the distilled variant based on Llama-8B with 8 billion parameters (Hugging Face name: deepseek-ai/ DeepSeek-R1-Distill-Llama-8B).
Computation Details. All experiments were conducted using the university’s high-performance computing (HPC) cluster managed via the Slurm workload manager. We used NVIDIA A100 (40 GB) and V100 (32 GB) GPUs for activation collection and intervention process. Each job was allocated 1 GPU, 1–10 CPU cores, and 50 GB of RAM.
Result Reliability. All quantitative results reported in the concept-erasure tables correspond to a single representative run. To verify the stability of our findings, we reran selected datasets (e.g., AG News) multiple times and observed identical outcomes and consistent relative rankings across methods. For instance, the top-performing method in the initial run remained the best in repeated trials. These consistent results indicate that the reported values are reliable and reproducible.
Package Implementation and Parameter Settings. Our implementation leverages several standard Python and deep learning libraries. The Hugging Face Transformers package was used to load and run language models (AutoTokenizer, AutoModelForCausalLM). The transformer_lens and sae_lens libraries were employed for model inspection and sparse autoencoder integration. The scikit-learn library was used for evaluation metrics such as ROC-AUC, while pandas and numpy supported data processing. Default configurations were used for all libraries, with reproducibility ensured by fixing random.seed(42).
Figure 7: JS distance vs. erasure ability for all partial erasure methods.
@@ Appendix E All SAE Neurons Analysis
While our salient neuron analysis provides valuable insight into the most strongly responding neurons, it considers only a narrow slice of the activation space (specifically, the top 80 neurons per concept). This limited scope may miss neurons that, although not highly ranked by mean activation, are still consistently active across multiple concepts and contribute to polysemanticity. To address this limitation, we broaden our analysis to include all neurons that exhibit non-zero activation for any concept, offering a more comprehensive view of concept overlap beyond the most salient neurons. This extended analysis is conducted only for SAEs, since in the dense base model, all neurons are active across all inputs, rendering such overlap statistics uninformative. The results are shown in Figure 6, which reports the intersection-over-union of active neurons across concepts. By capturing both highly active and more subtly engaged neurons, this analysis reveals a complete picture of polysemantic behavior. Consistent with our earlier findings shown in Figure 1, we observe that higher-capacity SAEs (e.g., 65k dimensions) exhibit lower neuron overlap than their lower-capacity counterparts. Together, Figures 1 and 6 demonstrate that while SAEs significantly reduce polysemanticity, they do not eliminate it entirely. Polysemantic neurons remain present, albeit to a lesser extent.
Appendix F Concept Erasure and JS Distance
As illustrated in Figure 7, all partial concept erasure methods exhibit a strong Pearson correlation with our JS separability score metric. Specifically, each method achieves a correlation coefficient greater than 0.7 with statistically significant $p$-values ($p<0.01$), indicating a robust relationship between concept separability and erasure precision. These results suggest that partial erasure techniques can effectively leverage the separability inherent in activation representations to suppress the targeted concept while minimizing interference with unrelated ones. Among the four partial erasure approaches evaluated, AURA and APP demonstrate the highest correlations (0.765 and 0.744, respectively), highlighting their superior ability to exploit distributional distinctions between concept activations. We attribute this performance advantage to the fact that both AURA and APP explicitly model the distributions of both target and auxiliary concepts as discussed in Subsection 4.4 (Comparative Analysis of APP and AURA).
Appendix G JS Separability: DeepSeek vs. Gemma
Table 7 presents the JS separability scores computed for both the Gemma-2-2B and DeepSeek-R1 models across all datasets. As indicated in the table, every SAE variant consistently achieves higher separability scores than its respective base model. This consistent improvement confirms that the incorporation of SAEs enhances the distinction between concept activation distributions, leading to more interpretable internal representations. Moreover, when comparing SAEs with the same capacity, those with higher sparsity (corresponding to lower $l_{0}$ values) exhibit greater separability. These results confirms that sparsity plays a critical role in improving the distinctness of concept distributions.
ID Primary Class Related Class Sentence (excerpt) Base 1621 SAE-16k 332 SAE-65k 5997
1 Business Science/Tech Microsoft, Dell big winners of Air Force deal … for software services worth \$500M 12.18-12.18 $20.15$ $61.67$
2 Business Sports Retailers bet on poker’s rising popularity … playing cards and chips … holiday shopping season $0.52$ $26.31$ $45.32$
3 Business World Intellectual property rights … US companies in China … WTO commitments 4.27-4.27 $12.16$ $79.13$
4 Sports World Doping cases hit record … 24 athletes ousted for drug-related violations … Olympics … no US team member flunked a test $0.64$ $20.24$ $49.52$
5 Sports Science/Tech Suggs to practice … Cleveland Browns running back Lee Suggs cleared by doctors to return … still some haziness about his medical condition $5.14$ $26.99$ $54.39$
6 World Science/Tech Ivan skirts Grand Cayman, flooding homes … winds near 250 km/h … uprooting trees, bursting canals, and flooding homes as it churned toward Cuba 12.35-12.35 $19.02$ $62.22$
Table 6: Example AG News sentences illustrating cross-class activation patterns. Each sentence’s primary class is shown alongside a semantically related secondary class. Activation values reflect the response of three separable neurons identified in Figure 2.
POS AG News Emotions DBpedia NER
Gemma
Base 0.343 0.366 0.322 0.405 0.308
SAE
(width_16k/l0_116)
0.539 0.650 0.431 0.621 0.581
SAE
(width_16k/l0_285)
0.425 0.567 0.389 0.543 0.511
SAE
(width_65k/l0_93)
0.600 0.709 0.446 0.680 0.621
SAE
(width_65k/l0_197)
0.549 0.626 0.407 0.644 0.537
DeepSeek
Base 0.367 0.366 0.291 0.436 0.325
SAE 0.597 0.698 0.643 0.718 0.698
Table 7: JS separability score comparison across datasets for Gemma and DeepSeek. Bold values indicate the best score per dataset.
Appendix H Concept Distribution Analysis
We analyze concept-conditional activation distributions on the AG News dataset, focusing on three neurons with high class separability in Figure 2: Base Neuron 1621, SAE-16k Neuron 332, and SAE-65k Neuron 5997.
To examine how these neurons encode secondary topical cues, consider the following example from the AG News dataset:
“Microsoft, Dell big winners of Air Force deal … for software services … worth \$500 million …”
The primary label of this sentence is Business, though it clearly involves elements of Science/Tech. The per-neuron activations for this sentence are:
  • Base Neuron 1621: 12.1797-12.1797
  • SAE-16k Neuron 332: $20.1485$
  • SAE-65k Neuron 5997: $61.6663$
Across all three neurons, the activations fall within the Business class density in Fig. 2 but are shifted toward the Science/Tech density. This pattern is consistent with polysemantic encoding; the neurons primarily represent the Business topic (capturing the correct class) while remaining sensitive to Science/Tech signals (e.g., “Microsoft,” “software”), which pulls the activation toward the Science/Tech mode without leaving the Business distribution.
Table 6 extends the analysis to six mixed-topic sentences. In every case, the three neurons produce activations that sit within the primary class distribution yet are shifted toward the related class. Absolute scales differ across neurons, so values should be interpreted relatively against each neuron’s class-conditional densities in Fig. 2 rather than compared across neurons. Overall, the pattern supports a polysemantic account; these neurons track the primary topic while remaining sensitive to secondary topical signals.